An experiment was performed to compare the fracture toughness of high-purity 18 Ni maraging steel with commercial-purity steel of the same type. For m = 30 specimens, the sample average toughness was x = 65.5 for the high-purity steel, whereas for n = 39 specimens of commercial steel y = 59.9. Because the high-purity steel is more expensive, its use for a certain application can be justified only if its fracture toughness exceeds that of commercial-purity steel by more than 5. Suppose that both toughness distributions are normal.

(a) Assuming that σ1 = 1.3 and σ2 = 1.1, test the relevant hypotheses using α = 0.001. (Use μ1μ2, where μ1 is the average toughness for high-purity steel and μ2 is the average toughness for commercial steel.)
State the relevant hypotheses.

H0: μ1μ2 = 5
Ha: μ1μ2 ≤ 5H0: μ1μ2 = 5
Ha: μ1μ2 > 5    H0: μ1μ2 = 5
Ha: μ1μ2 < 5H0: μ1μ2 = 5
Ha: μ1μ2 ≠ 5


Calculate the test statistic and determine the P-value. (Round your test statistic to two decimal places and your P-value to four decimal places.)

z =
P-value =



State the conclusion in the problem context.

Fail to reject H0. The data suggests that the fracture toughness of high-purity steel exceeds that of commercial-purity steel by more than 5.Reject H0. The data suggests that the fracture toughness of high-purity steel exceeds that of commercial-purity steel by more than 5.     Reject H0. The data does not suggest that the fracture toughness of high-purity steel exceeds that of commercial-purity steel by more than 5.Fail to reject H0. The data does not suggest that the fracture toughness of high-purity steel exceeds that of commercial-purity steel by more than 5.


(b) Compute β for the test conducted in part (a) when μ1μ2 = 6. (Round your answer to four decimal places.)
Answer #1

a)

H0: μ1μ2 = 5
Ha: μ1μ2 > 5

b)

pop 1 pop 2
sample mean x = 65.50 59.90
std deviation σ= 1.300 1.100
sample size n= 30 39
std error σx1-x2=√(σ21/n122/n2)    = 0.296
test stat z =(x1-x2o)/σx1-x2     = 2.03
p value : = 0.0212 (from excel:1*normsdist(-2.03)

Fail to reject H0. The data does not suggest that the fracture toughness of high-purity steel exceeds that of commercial-purity steel by more than 5.

b_)

standard error=(√(σ12/n1+σ12/n1))= 0.2956
for 0.001 level and right tail critival Zα= 3.090
rejecf Ho if x>= Δo +Zα*σx or x>= 5.9133
P(Type II error) =P(Xbar<5.913| Δ=6)=P(Z<(5.9133-6)/0.296)=P(Z<-0.29)=0.3859